OpenStudy (anonymous):

My last question. This one is kind of hard to work out on a computer, so I understand if you don't want to answer it. Just typing out what to do first is alright, no need for drawings (unless you want to) Factor x^4+4x^3-52x^2+20x+75=0 Root given: -4+ (root) 31 It's supposed to be synthetic division.

OpenStudy (campbell_st):

well just looking at it and a quick calculation (x + 1) is a factor since f(-1) = 0 the given root just indicates to me that 1 factor is a quadratic. so using the general quadratic formula you know that \[x = -4 + \sqrt{31}\] then \[x + 4 = \sqrt{31}\] so \[(x + 4)^2 = 31\] by expanding and subtracting 31 you should be able to get the quadratic factor. all you need now is the other linear factor. hope this helps.

OpenStudy (anonymous):

Thanks a lot for your help! I'm now done with my math packet :) But I still have to study more for my final tomorrow.

OpenStudy (campbell_st):

good luck

OpenStudy (anonymous):

Thanks ^^